# Rational powers, invariant ideals, and the summation formula

### v2

@@ 4. Rees package and the summation formula
Write $\Gamma_{1}=\mathscr{Q}_{1}+\mathbb{R}_{\geqslant 0}^{n_{1}}$ and $\Gamma_{2}=\mathscr{Q}_{2}+\mathbb{R}_{\geqslant 0}^{n_{2}}$ where $\mathscr{Q}_{1}$ and $\mathscr{Q}_{2}$ are rational polytopes not containing the origin. By [46, Proposition 2.3] the second statement holds for ${\operatorname{conv}}(\mathscr{Q}_{1},\mathscr{Q}_{2})$. By [8, Proposition 1.31] and its proof we have ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})={\operatorname{conv}}(\mathscr{Q}_{1},\mathscr{Q}_{2})+\mathbb{R}_{\geqslant 0}^{n_{1}+n_{2}}$, and from this both statements follow for ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})$. ∎
@@
Consider the polynomial rings $R=\mathbb{Q}[x]$ and $S=\mathbb{Q}[y]$, and the prime ideals $I=(x^{2}+1)\subset R$, $J=(y^{2}+1)\subset S$. Since $\overline{R}\left[\dfrac{x^{2}+1}{x^{2}+1}\right]=\overline{R}=R$ and $IR$ has only one minimal prime we obtain $\left|{\operatorname{\mathscr{RV}}}(I)\right|=1$. Similarly $\left|{\operatorname{\mathscr{RV}}}(J)\right|=1$. Using Macaulay2 [27] we conclude that the ideal $IT+JT=(x^{2}+1,y^{2}+1)$ is integrally closed and that it has two minimal primes. Therefore $|\mathcal{R}\mathcal{V}(I+J)|\geqslant 2$ by [35, Discussion 10.1.3].
5. Asymptotic rational powers over the complex numbers
@@
Let $R$ and $S$ be $\mathbb{C}$-algebras, and $I\subset R$ and $J\subset S$ nonzero ideals. We denote by $IT+JT$ the $T$-ideal $IT+JT$, where $T:=R\otimes_{\mathbb{C}}S$.
@@
Adopt Section missing 5 with $R$ and $S$ being finitely generated $\mathbb{C}$-algebras and normal domains. There exists a $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
$\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{I^{\alpha}}\,\overline{J^{w-\alpha}}\,T\subseteq\overline{\left(IT+JT\right)^{w}}\subseteq\overline{I^{\tau}}T+\overline{J^{w-\tau}}T=\overline{(IT)^{\tau}}+\overline{(JT)^{w-\tau}}.$
@@
We note that by [47, Lemma 1.2] (see also [41, Lemma 9.5.31]) we have
@@
Before presenting the proof of the theorem, we need some preparatory results. We start with the following statement, which is a generalization of [34, Theorem 2.3] that is suitable for our purposes. First, we need the following definitions.
@@
We only include relevant details of how a modification of the proof of [34, Theorem 2.3] leads to our statement. We refer the reader to the proof of [34, Theorem 2.3] for any undefined notation, which we will follow closely.
@@
We proceed with the reverse inclusion. Let $H$ be a non-coordinate supporting hyperplane of ${\operatorname{NP}}(I)$ defined by $\langle\mathbf{h},\mathbf{X}\rangle=h$ where $\mathbf{h}=(h_{1},\ldots,h_{d})\in{\mathbb{Z}}_{\geqslant 0}^{d}$, $h\in\mathbb{Z}_{>0}$, and $\mathbf{X}=(X_{1},\ldots,X_{d})$. Let $I_{H}$ be the ideal generated by the monomials $\mathbf{x}^{\mathbf{e}}$ with $\langle\mathbf{h},\mathbf{e}\rangle\geqslant h$. It suffices to show that the conclusion of the theorem holds for $I_{H}$. Let $\mathfrak{p}=(Y_{1},\ldots,Y_{t})R^{\prime}$. By Section missing 2 and [35, Corollary 5.4.2], the rational powers of $\mathfrak{p}^{h}R^{\prime}_{\mathfrak{q}}$ have no embedded primes for every prime ideal $\mathfrak{p}\subseteq\mathfrak{q}$. Thus, it suffices to show that $\overline{\left(\mathfrak{p}^{h}R^{\prime}_{\mathfrak{p}}\right)^{w}}=\mathfrak{p}^{\lceil wh\rceil}R^{\prime}_{\mathfrak{p}}$, which follows from $\mathfrak{p}R^{\prime}_{\mathfrak{p}}$ being a prime ideal generated by a regular sequence, as in this case the powers $\mathfrak{p}^{n}R^{\prime}_{\mathfrak{p}}$ are integrally closed and $\operatorname{gr}_{R^{\prime}_{\mathfrak{p}}}(\mathfrak{p}R^{\prime}_{\mathfrak{p}}):=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\mathfrak{p}^{n}R^{\prime}_{\mathfrak{p}}/\mathfrak{p}^{n+1}R^{\prime}_{\mathfrak{p}}$ is a domain. ∎
@@
Let $(V_{1},\eta_{1})$ and $(V_{2},\eta_{2})$ be two divisorial valuation rings containing the field of complex numbers and such that $V:=V_{1}\otimes_{\mathbb{C}}V_{2}$ is Noetherian. Let $I=(\eta_{1}^{n_{1}})$ and $J=(\eta_{2}^{n_{2}})$ be ideals of $V_{1}$ and $V_{2}$, respectively. For every $\tau,w\in\mathbb{Q}_{\geqslant 0}$ with $\tau\leqslant w$, we have the following inclusion of $V$-ideals
@@
By [56, Theorem 6(e)] $\eta_{1},\eta_{2}$ is a grsop of $V$ (see Section missing 5). Thus by Theorem 5.2 the ideal $\overline{\left(IT+JT\right)^{w}}$ is generated by the monomials $\eta_{1}^{e_{1}}\eta_{2}^{e_{2}}$ with $n_{2}e_{1}+n_{1}e_{2}\geqslant wn_{1}n_{2}$. Therefore, for any such $\eta_{1}^{e_{1}}\eta_{2}^{e_{2}}$ we must have $e_{1}\geqslant\tau n_{1}$ or $e_{2}\geqslant(w-\tau)n_{2}$. The result follows as $\overline{I^{\tau}}$ is generated by $\eta_{1}^{\lceil\tau n_{1}\rceil}$ and $\overline{J^{w-\tau}}$ by $\eta_{2}^{\lceil(w-\tau)n_{2}\rceil}$. ∎
The following lemma offers a vanishing result for rational powers of exceptional divisors that can be of independent utility and interest.
@@
Adopt Section missing 2. Let $v:X^{+}\longrightarrow X$ be the normalization of the blowup of $X$ along $\mathcal{I}$, and let $E$ be its exceptional divisor. Then there exists a positive integer $n$ such that if $\beta\in\mathbb{Q}$ and $\beta\geqslant n$ we have $R^{1}v_{*}\left(\mathcal{O}_{X^{+}}(-\lceil\beta E\rceil)\right)=0$.
@@
Set $\mathscr{T}:=\mathcal{O}_{X^{+}}(-E)$. By Section missing 2 we have $\overline{\mathscr{T}^{\beta}}=\mathcal{O}_{X^{+}}\left(-\lceil\beta E\rceil\right)$ for every $\beta\in\mathbb{Q}_{\geqslant 0}$. Notice that there exists $m\in\mathbb{Z}_{>0}$ such that for every $\beta\geqslant m$ we have $\overline{\left(\mathscr{T}^{+}\right)^{\beta+m}}=\overline{\left(\mathscr{T}^{+}\right)^{\beta}}\,\,\overline{\left(\mathscr{T}^{+}\right)^{m}}$. Indeed, for any finite affine cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $X^{+}$, this holds locally on each $U_{\lambda}$ by Section missing 2(3), producing finitely many $m_{\lambda}$. Thus we take can take $m:={\operatorname{lcm}}(m_{\lambda}\mid\lambda\in\Lambda)$. Therefore for every $\beta\geqslant m$ we have
@@
where the last isomorphism holds by flatness. Since $m\leqslant\beta-m\lfloor\beta/m\rfloor+m\leqslant 2m$, by Section missing 2 there are finitely many distinct ideal sheaves of the form $\overline{\left(\mathscr{T}^{+}\right)^{\beta-m\lfloor\beta/m\rfloor+m}}$. Since $\mathcal{O}_{X^{+}}(-E)$ is ample for $v$ [54, tag 01OF, Lemma 31.32.13], by Serre’s vanishing theorem [41, Theorem 1.7.6] there exists $n\in\mathbb{Z}_{>0}$ such that for every $\beta\geqslant n$
@@
proving the first inclusion. Moreover, also by Section missing 2, the last equality holds. We proceed with the second inclusion to show that there exists $w_{0}\in\mathbb{Q}_{>0}$ such that
@@
for $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$. The last equality follows from Künneth formula (see [54, tag 0BEC, Lemma 33.29.1] and [31, Proposition III.8.5]) using Theorem 2.2 and the facts that $v_{1*}\left(\mathcal{O}_{X_{1}^{+}}\right)=\mathcal{O}_{X_{1}}$ and $v_{2*}\left(\mathcal{O}_{X_{2}^{+}}\right)=\mathcal{O}_{X_{2}}$ [31, proof of Corollary III.11.4]. Thus, (missing) 6 will follow by taking global sections, see Section missing 2.
We proceed to prove Claim 1, for which we can restrict to a product of affine schemes $U_{1}\times U_{2}\subset Y^{+}$, where $U_{1}={\operatorname{Spec\,}}(A_{1})\subset X_{1}^{+}$, $V={\operatorname{Spec\,}}(A_{2})\subset X_{2}^{+}$, and $A_{1}$ and $A_{2}$ are Noetherian normal domains. Therefore $\mathcal{I}^{+}$ and $\mathcal{T}^{+}$ can be identified with principal ideals $(a_{1})\subseteq A_{1}$ and $(a_{2})\subseteq A_{2}$, respectively, and $q_{1}^{-1}\mathcal{I}^{+}+q_{2}^{-1}\mathcal{T}^{+}$ with $(a_{1})A+(a_{2})A$ where $A:=A_{1}\otimes_{\mathbb{C}}A_{2}$. By Section missing 2 and [35, Corollary 5.4.2] the associated primes of the rational powers $\overline{(a_{1})^{\tau}}$ and $\overline{(a_{2})^{w-\tau}}$ all have codimension one. By [30, Corollary 2.8] the associated primes of $(a_{1})A+(a_{2})A$ are sums $\mathfrak{p}_{1}A+\mathfrak{p}_{2}A$ with $\mathfrak{p}_{i}$ associated prime of $(a_{i})$, $i=1,2$. Moreover, rational powers localize by Section missing 2. Thus, to show $\overline{\left((a_{1})A+(a_{2})A\right)^{w}}\subseteq\overline{(a_{1})^{\tau}}A+\overline{(a_{2})^{w-\tau}}A$, it suffices to do so in the localizations $(A_{1})_{\mathfrak{p}_{1}}\otimes_{\mathbb{C}}(A_{2})_{\mathfrak{p}_{2}}$. But the latter follows from Section missing 5, finishing the proof of Claim 1.
@@
$R^{1}v_{1*}\left(\overline{\left(\mathcal{I}^{+}\right)^{\beta}}\right)=0\qquad\text{and}\qquad R^{1}v_{2*}\left(\overline{\left(\mathcal{T}^{+}\right)^{\beta}}\right)=0.$
@@
By working locally in affine open subsets, Theorem 5.1 implies the following more general statement (cf. [47, Theorem 0.3]). As in the proof of Theorem 5.1, we denote by $p^{-1}\mathcal{J}$, resp. $q^{-1}\mathcal{J}$, the ideal sheaves $p^{-1}\mathcal{J}\cdot\mathcal{O}_{X\times Y}$, resp. $q^{-1}\mathcal{J}\cdot\mathcal{O}_{X\times Y}$.
@@
Let $X,Y$ be normal complex varieties and $\mathcal{I}\subseteq\mathcal{O}_{X},\mathcal{T}\subseteq\mathcal{O}_{y}$ nonzero ideal sheaves. Let $p:X\times Y\longrightarrow X$ and $q:X\times Y\longrightarrow Y$ be the canonical projections. There exists a $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
$\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}p^{-1}\overline{\mathcal{I}^{\alpha}}\,\,q^{-1}\overline{\mathcal{T}^{w-\alpha}}\subseteq\overline{\left(p^{-1}\mathcal{I}\,+\,q^{-1}\mathcal{T}\right)^{w}}\subseteq\overline{\left(p^{-1}\mathcal{I}\right)^{\tau}}\,+\,\overline{\left(q^{-1}\mathcal{T}\right)^{w-\tau}}.$
Similarly as in [47, Proposition 2.1], the formula in Theorem 5.1 can be used to approximate rational powers up to $\mathfrak{m}$-adic neighborhoods. This is also related to the main theorem in [19, 20].
@@
Adopt Section missing 5 with $R$ a finitely generated $\mathbb{C}$-algebra and normal domain, and $S$ a polynomial ring with homogeneous maximal ideal $\mathfrak{m}$. For every $n\in{\mathbb{Z}}_{\geqslant 0}$, there exists a $w_{0}=w_{0}(n)\in\mathbb{Q}_{>0}$ such that if $w,\ell\geqslant w_{0}$ then
@@
The first statement follows from Theorem 5.1 and since the powers of $\mathfrak{m}$ are integrally closed, then $\overline{\left(\mathfrak{m}^{n}\right)^{\ell}}=\mathfrak{m}^{\lceil n\ell\rceil}$. The second statement also follows from Theorem 5.1 and by the inclusions $\overline{J^{\ell}}\subseteq\overline{J^{\lfloor\ell\rfloor}}\subseteq J^{\lfloor\ell\rfloor-\dim(S)+1}$, which hold by the Briançon-Skoda Theorem [45] (see also [35, Corollary 13.3.4]). ∎

### v3

@@ 4. Rees package and the summation formula
Write $\Gamma_{1}=\mathscr{Q}_{1}+\mathbb{R}_{\geqslant 0}^{n_{1}}$ and $\Gamma_{2}=\mathscr{Q}_{2}+\mathbb{R}_{\geqslant 0}^{n_{2}}$ where $\mathscr{Q}_{1}$ and $\mathscr{Q}_{2}$ are rational polytopes not containing the origin. By [48, Proposition 2.3] the second statement holds for ${\operatorname{conv}}(\mathscr{Q}_{1},\mathscr{Q}_{2})$. By [9, Proposition 1.31] and its proof we have ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})={\operatorname{conv}}(\mathscr{Q}_{1},\mathscr{Q}_{2})+\mathbb{R}_{\geqslant 0}^{n_{1}+n_{2}}$, and from this both statements follow for ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})$. ∎
@@
Consider the polynomial rings $R=\mathbb{Q}[x]$ and $S=\mathbb{Q}[y]$, and the prime ideals $I=(x^{2}+1)\subset R$, $J=(y^{2}+1)\subset S$. Since $\overline{R}\left[\dfrac{x^{2}+1}{x^{2}+1}\right]=\overline{R}=R$ and $IR$ has only one minimal prime we obtain $\left|{\operatorname{\mathscr{RV}}}(I)\right|=1$. Similarly $\left|{\operatorname{\mathscr{RV}}}(J)\right|=1$. Using Macaulay2 [29] we conclude that the ideal $IT+JT=(x^{2}+1,y^{2}+1)$ is integrally closed and that it has two minimal primes. Therefore $|\mathcal{R}\mathcal{V}(I+J)|\geqslant 2$ by [38, Discussion 10.1.3].
5. Asymptotic rational powers over algebraically closed fields
@@
Let $\mathbb{k}$ be an algebraically closed field. Let $R$ and $S$ be $\mathbb{k}$-algebras, and $I\subset R$ and $J\subset S$ nonzero ideals. We denote by $IT+JT$ the $T$-ideal $IT+JT$, where $T:=R\otimes_{\mathbb{k}}S$.
@@
Adopt Section missing 5 with $R$ and $S$ being finitely generated $\mathbb{k}$-algebras and normal domains. There exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
@@
We note that by [49, proof of Lemma 1.2] we have
@@
Before presenting the proof of the theorem, we need some preparatory results. We start with the following statement, which is a generalization of [37, Theorem 2.3] that is suitable for our purposes. First, we need the following definitions.
@@
We only include relevant details of how a modification of the proof of [37, Theorem 2.3] leads to our statement. We refer the reader to the proof of [37, Theorem 2.3] for any undefined notation, which we will follow closely.
@@
We proceed with the reverse inclusion. Let $H$ be a non-coordinate supporting hyperplane of ${\operatorname{NP}}(I)$ defined by $\langle\mathbf{h},\mathbf{X}\rangle=h$ where $\mathbf{h}=(h_{1},\ldots,h_{d})\in{\mathbb{Z}}_{\geqslant 0}^{d}$, $h\in\mathbb{Z}_{>0}$, and $\mathbf{X}=(X_{1},\ldots,X_{d})$. Let $I_{H}$ be the ideal generated by the monomials $\mathbf{x}^{\mathbf{e}}$ with $\langle\mathbf{h},\mathbf{e}\rangle\geqslant h$. It suffices to show that the conclusion of the theorem holds for $I_{H}$. Let $\mathfrak{p}=(Y_{1},\ldots,Y_{t})R^{\prime}$. By Section missing 2 and [38, Corollary 5.4.2], the rational powers of $\mathfrak{p}^{h}R^{\prime}_{\mathfrak{q}}$ have no embedded primes for every prime ideal $\mathfrak{p}\subseteq\mathfrak{q}$. Thus, it suffices to show that $\overline{\left(\mathfrak{p}^{h}R^{\prime}_{\mathfrak{p}}\right)^{w}}=\mathfrak{p}^{\lceil wh\rceil}R^{\prime}_{\mathfrak{p}}$, which follows from $\mathfrak{p}R^{\prime}_{\mathfrak{p}}$ being a prime ideal generated by a regular sequence, as in this case the powers $\mathfrak{p}^{n}R^{\prime}_{\mathfrak{p}}$ are integrally closed and $\operatorname{gr}_{R^{\prime}_{\mathfrak{p}}}(\mathfrak{p}R^{\prime}_{\mathfrak{p}}):=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\mathfrak{p}^{n}R^{\prime}_{\mathfrak{p}}/\mathfrak{p}^{n+1}R^{\prime}_{\mathfrak{p}}$ is a domain. ∎
@@
Let $\mathbb{k}$ be an algebraically closed field. Let $(V_{1},\eta_{1})$ and $(V_{2},\eta_{2})$ be two divisorial valuation rings containing $\mathbb{k}$ and such that $V:=V_{1}\otimes_{\mathbb{k}}V_{2}$ is Noetherian. Let $I=(\eta_{1}^{n_{1}})$ and $J=(\eta_{2}^{n_{2}})$ be ideals of $V_{1}$ and $V_{2}$, respectively. For every $\tau,w\in\mathbb{Q}_{\geqslant 0}$ with $\tau\leqslant w$, we have the following inclusion of $V$-ideals
@@
By [57, Theorem 6(e)] and [5, Chapter V, §17.2, Corollary], $V$ is a regular domain and $\eta_{1},\eta_{2}$ is a grsop of $V$ (see Section missing 5). Thus by Theorem 5.2 the ideal $\overline{\left(IT+JT\right)^{w}}$ is generated by the monomials $\eta_{1}^{e_{1}}\eta_{2}^{e_{2}}$ with $n_{2}e_{1}+n_{1}e_{2}\geqslant wn_{1}n_{2}$. Therefore, for any such $\eta_{1}^{e_{1}}\eta_{2}^{e_{2}}$ we must have $e_{1}\geqslant\tau n_{1}$ or $e_{2}\geqslant(w-\tau)n_{2}$. The result follows as $\overline{I^{\tau}}$ is generated by $\eta_{1}^{\lceil\tau n_{1}\rceil}$ and $\overline{J^{w-\tau}}$ by $\eta_{2}^{\lceil(w-\tau)n_{2}\rceil}$. ∎
The following lemma offers a vanishing result for rational powers of exceptional divisors of normalized blowups that can be of independent utility and interest.
@@
Adopt Section missing 2. Let $v:X^{+}\longrightarrow X$ be the normalization of the blowup of $X$ along $\mathcal{I}$, and let $E$ be its exceptional divisor, i.e., $\mathcal{I}\cdot\mathcal{O}_{X^{+}}=\mathcal{O}_{X^{+}}(-E)$. Then there exists a positive integer $n$ such that if $\beta\in\mathbb{Q}$ and $\beta\geqslant n$ we have $R^{1}v_{*}\left(\mathcal{O}_{X^{+}}(-\lceil\beta E\rceil)\right)=0$.
@@
Set $\mathscr{T}:=\mathcal{O}_{X^{+}}(-E)$. By Section missing 2 we have $\overline{\mathscr{T}^{\beta}}=\mathcal{O}_{X^{+}}\left(-\lceil\beta E\rceil\right)$ for every $\beta\in\mathbb{Q}_{\geqslant 0}$. Notice that there exists $m\in\mathbb{Z}_{>0}$ such that for every $\beta\geqslant m$ we have $\overline{\left(\mathscr{T}^{+}\right)^{\beta+m}}=\overline{\left(\mathscr{T}^{+}\right)^{\beta}}\,\,\overline{\left(\mathscr{T}^{+}\right)^{m}}$. Indeed, for any finite affine cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $X^{+}$, this holds locally on each $U_{\lambda}$ by Section missing 2(3), producing finitely many $m_{\lambda}$. So, we take can take $m:={\operatorname{lcm}}(m_{\lambda}\mid\lambda\in\Lambda)$. Therefore for every $\beta\geqslant m$ we have
@@
where the last isomorphism holds by flatness. Since $m\leqslant\beta-m\lfloor\beta/m\rfloor+m\leqslant 2m$, by Section missing 2 there are finitely many distinct ideal sheaves of the form $\overline{\left(\mathscr{T}^{+}\right)^{\beta-m\lfloor\beta/m\rfloor+m}}$. Since $\mathcal{O}_{X^{+}}(-E)$ is ample, by Serre’s vanishing theorem [33, Proposition III.5.3] there exists $n\in\mathbb{Z}_{>0}$ such that for every $\beta\geqslant n$
@@
proving the first inclusion. We proceed with the second inclusion to show that there exists $w_{0}\in\mathbb{Q}_{>0}$ such that
@@
for $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$. The last equality follows from Künneth formula (see [55, tag 0BEC, Lemma 33.29.1] and [34, Proposition III.8.5]) using Theorem 2.2 and the facts that $v_{1*}\left(\mathcal{O}_{X_{1}^{+}}\right)=\mathcal{O}_{X_{1}}$ and $v_{2*}\left(\mathcal{O}_{X_{2}^{+}}\right)=\mathcal{O}_{X_{2}}$ [34, proof of Corollary III.11.4]. Thus, (missing) 6 will follow by taking global sections, see Section missing 2.
We proceed to prove Claim 1, for which we can restrict to a product of affine schemes $U_{1}\times U_{2}\subset Y^{+}$, where $U_{1}={\operatorname{Spec\,}}(A_{1})\subset X_{1}^{+}$, $V={\operatorname{Spec\,}}(A_{2})\subset X_{2}^{+}$, and $A_{1}$ and $A_{2}$ are Noetherian normal domains. Therefore $\mathcal{I}^{+}$ and $\mathcal{T}^{+}$ can be identified with principal ideals $(a_{1})\subseteq A_{1}$ and $(a_{2})\subseteq A_{2}$, respectively, and $q_{1}^{-1}\mathcal{I}^{+}+q_{2}^{-1}\mathcal{T}^{+}$ with $(a_{1})A+(a_{2})A$ where $A:=A_{1}\otimes_{\mathbb{k}}A_{2}$. By Section missing 2 and [38, Corollary 5.4.2] the associated primes of the rational powers $\overline{(a_{1})^{\tau}}$ and $\overline{(a_{2})^{w-\tau}}$ all have codimension one. By [32, Corollary 2.8] the associated primes of $(a_{1})A+(a_{2})A$ are sums $\mathfrak{p}_{1}A+\mathfrak{p}_{2}A$ with $\mathfrak{p}_{i}$ associated prime of $(a_{i})$, $i=1,2$. Moreover, rational powers localize by Section missing 2. Thus, to show $\overline{\left((a_{1})A+(a_{2})A\right)^{w}}\subseteq\overline{(a_{1})^{\tau}}A+\overline{(a_{2})^{w-\tau}}A$, it suffices to do so in the localizations $(A_{1})_{\mathfrak{p}_{1}}\otimes_{\mathbb{k}}(A_{2})_{\mathfrak{p}_{2}}$. But the latter follows from Section missing 5, finishing the proof of Claim 1.
@@
(7) $R^{1}v_{1*}\left(\overline{\left(\mathcal{I}^{+}\right)^{\beta}}\right)=0\qquad\text{and}\qquad R^{1}v_{2*}\left(\overline{\left(\mathcal{T}^{+}\right)^{\beta}}\right)=0.$
@@
We note that if for certain ideals $I$ and $J$ the vanishings in (7) hold for every $\beta\geqslant 0$, then the proof of Theorem 5.1 shows that these ideals satisfy the conclusion of Section missing 1.
By working locally in affine open subsets, Theorem 5.1 implies the following more general statement (cf. [49, Theorem 0.3]). As in the proof of Theorem 5.1, we denote by $p^{-1}\mathcal{J}$, resp. $q^{-1}\mathcal{J}$, the ideal sheaves $p^{-1}\mathcal{J}\cdot\mathcal{O}_{X\times Y}$, resp. $q^{-1}\mathcal{J}\cdot\mathcal{O}_{X\times Y}$.
@@
Let $X,Y$ be normal varieties over an algebraically closed field $\mathbb{k}$. Let $\mathcal{I}\subseteq\mathcal{O}_{X},\mathcal{T}\subseteq\mathcal{O}_{y}$ be nonzero ideal sheaves, and $p:X\times Y\longrightarrow X$, $q:X\times Y\longrightarrow Y$ the canonical projections. There exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
$\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}p^{-1}\overline{\mathcal{I}^{\alpha}}\,\,q^{-1}\overline{\mathcal{T}^{w-\alpha}}\subseteq\overline{\left(p^{-1}\mathcal{I}\,+\,q^{-1}\mathcal{T}\right)^{w}}\subseteq p^{-1}\overline{\left(\mathcal{I}\right)^{\tau}}\,+\,q^{-1}\overline{\left(\mathcal{T}\right)^{w-\tau}}.$
Similarly as in [49, Proposition 2.1], the formula in Theorem 5.1 can be used to approximate rational powers up to $\mathfrak{m}$-adic neighborhoods. This is also related to the main theorem in [23, 24].
[…]
Adopt Section missing 5 with $R$ a finitely generated $\mathbb{k}$-algebra and normal domain, and $S$ a polynomial ring with homogeneous maximal ideal $\mathfrak{m}$. For every $n\in{\mathbb{Z}}_{\geqslant 0}$, there exists $w_{0}=w_{0}(n)\in\mathbb{Q}_{>0}$ such that if $w,\ell\geqslant w_{0}$ then
@@
The first statement follows from Theorem 5.1 and since the powers of $\mathfrak{m}$ are integrally closed, then $\overline{\left(\mathfrak{m}^{n}\right)^{\ell}}=\mathfrak{m}^{\lceil n\ell\rceil}$. The second statement also follows from Theorem 5.1 and by the inclusions $\overline{J^{\ell}}\subseteq\overline{J^{\lfloor\ell\rfloor}}\subseteq J^{\lfloor\ell\rfloor-\dim(S)+1}$, which hold by the Briançon-Skoda Theorem [47] (see also [38, Corollary 13.3.4]). ∎